# Towards Human-AI Deliberation: Design and Evaluation of LLM-Empowered Deliberative AI for AI-Assisted Decision-Making

### v1

@@ 3.1. Unpacking Decision-Making
In this work, we dissect a decision-making problem and human-AI thoughts into four structured components (Figure 2 (a)). (1) Decision means the overall decision to make for a specific problem. (2) Dimension refers to an aspect that people consider when reaching an overall decision. For instance, in a graduate admission task, dimensions could be applicants’ academic excellence, research ability, etc. Taking tabular task data (commonly used in decision-making (Ghai et al., 2021; Wang and Yin, 2021)) as an example, a dimension may be an attribute or a set of related attributes. (3) Opinion on a dimension denotes one’s views on the impact of a dimension on the overall decision (e.g., an applicant’s academic excellence contributes +50% to the overall chance of being admitted). (4) Evidence is the basis of an opinion. For humans, evidence can encompass facts, heuristics, past experiences, personal beliefs, interpretations, and even creative thoughts (Simon, 1990, 1997). In contrast, AI’s evidence primarily originates from information and knowledge embedded in the training data.
3.1.1. When and Where to Deliberate?
In the scope of this paper, we propose to conduct Human-AI Deliberation in “Weigh the Evidence” (Step 4 in the seven decision-making steps). For one thing, disagreements between human’s thoughts and AI’s suggestions often surface when they assess alternatives in this step. Diverse interpretations of evidence in Step 4 are likely to lead to conflicting opinions and disparate outcomes in steps 5-7 (Lai et al., 2021). For another, in existing AI-assisted decision-making tasks (Lai et al., 2021; Lai and Tan, 2019; Bansal et al., 2021), steps 1-3 are usually settled in advance and there is no need for much deliberation. We propose to have humans and AI deliberate their opinions at the detailed Dimension level (i.e., to what extent a dimension supports or opposes the final decision) rather than merely talking about accepting/rejecting one’s overall decision. In this process, both human and AI need to support their arguments with evidence and weigh the presented evidence according to credibility and probative value (Baumann et al., 2014).
3.1.2. What to Deliberate on?
In real-world practices, a deliberation process in Step 4 is often informed by the importance of evidence interpreted in a particular context (Cartwright and Stegenga, 2011). In information theory, such “importance” is quantified as the (probabilistic) weight of evidence (WoE) for a given hypothesis (and against alternative hypotheses) (Peirce, 2014; Good, 1950; Gough, 2007). WoE is well studied and widely adopted in various decision-making domains (Alvarez-Melis et al., 2021). This is because its literal meaning is easy to understand by lay users and its probabilistic calculation is simple to implement in practical applications (Saaty, 2008; Gough, 2007).
WoE is the core content of deliberation in our proposed framework. It corresponds to “opinion” – the importance of a dimension to a final decision given one’s interpretation of evidence. In the remainder of this paper, we use the term opinion (as well as its synonyms e.g., viewpoint) and WoE interchangeably.
It is crucial to note that during evidence weighing, both humans and AI possess strengths and limitations. For instance, humans may offer novel insights not present in AI training data, yet they can be constrained by time and cognitive biases, as per the “Bounded Rationality Model” (Simon, 1990, 1997). AI excels in data-driven knowledge and computation, but is known to be susceptible to biases in training data and falters in the face of out-of-distribution cases (Liu et al., 2021). These inherent human and AI shortcomings can result in potentially incorrect, incomplete, or biased evidence weighting, leading to opinion discrepancy and even decision errors. Hence, fostering humans’ deliberation on conflicting evidence weights could be advantageous.
@@ 3.1.2. What to Deliberate on?
Figure 2. The framework for Human-AI Deliberation. (A) Illustrates the Weight of Evidence (WoE) concept in decision-making, showcasing how decision-makers assess evidence across dimensions to shape opinions and arrive at a final decision. (B) Presents the Architecture for Human-AI Deliberation, with key activities (shown in grey boxes) and potential design space (shown in dashed-line boxes).
@@
3.2. Human-AI Deliberation: A Framework
Deliberation is beneficial for enhancing decision-making by integrating diverse perspectives, improving the quality of solutions by considering various viewpoints, fostering critical thinking, and increasing the likelihood of reaching informed and well-thought-out conclusions (Gracia, 2003). Although deliberation is commonly used in democracy and political settings (Black et al., 2010; Habermas, 2005; Thompson, 2008), its idea can be applied to other decision-making tasks such as clinical diagnosis (Schaekermann et al., 2019; Inguaggiato et al., 2019; Preisz, 2019), criminal justice (Devine et al., 2001; Van Berkel et al., 2019), etc. Building upon the WoE decision-making approach, we propose Human-AI Deliberation, a framework to stimulate deliberative processes involving both humans and AI (Figure 2). This framework comprises the following essential activities:
  • Elicitation of Thoughts: Human and AI start with articulating their dimension-level perspectives on the decision problem. While AI presenting its “thoughts” (e.g., in the form of feature importance explanation) is rather common in AI-assisted decision-making (Liao and Varshney, 2021), this activity also encourages individuals to clarify their ideas and examine their reasoning, which prompts analytical thinking in human (Milkman et al., 2009; Buçinca et al., 2021). Two aspects of this activity require careful design. First, AI thought elicitation demands a good balance between human information needs and interpretability (Mishra and Rzeszotarski, 2021; Abdul et al., 2020). Prior research suggested that certain mechanisms AI applies to make inferences and predictions, although informative, are not comprehensible to humans (Miller, 2019). Hence, the design of AI’s “thought” explanations should be critical for this activity. Second, human thought elicitation, while encouraging thoughtful reasoning (Milkman et al., 2009; Buçinca et al., 2021), can impose a potential workload. It thus demands suitable, friendly interface designs.
  • Alignment of Human-AI Thoughts: As human’s and AI’s viewpoints and the process they form those viewpoints may diverge (Kaufman and Kirsh, 2022; Miller, 2019; Holzinger et al., 2019), this activity is tasked with establishing a common language for the two parties to compare their WoE and determine the extent of discrepancy, if any. Proper assessment and presentation of human-AI WoE differences can help effectively navigate humans’ attention and efforts in the subsequent activities (Boggust et al., 2022).
  • Discussion: This activity features constructive discussions that enable humans and AI to substantiate their opinions via communication, clarifying choices of evidence, and the basis for assigned weights. It has the potential benefits of cultivating critical thinking, mitigating biases, and raising awareness of differences between multiple parties (Milkman et al., 2009; Larson et al., 1994; Hirokawa, 1985). This activity has a huge design space and usually needs to be tailored to specific decision tasks. It entails considering various discussion-related factors, including content, style, leadership (who initiates and leads the conversation), duration, and so on. A typical solution is transferring human-human discussion practices (Milkman et al., 2009; Larson et al., 1994; Hirokawa, 1985) to the human-AI discussion contexts.
@@ 3.2. Human-AI Deliberation: A Framework
In summary, the proposed Human-AI Deliberation framework consists of four interlinked activities and requires proper designs (the potential design space is illustrated in the dashed box in Figure 2 (b)) of a decision-making interface and an AI with the ability to deliberate with humans. Since the interface design is task-dependent, we only introduce the design of Deliberative AI in the next subsection. Notably, not all activities in this framework are necessarily required for Human-AI Deliberation. Furthermore, we do not intend to exhaust all possible forms of Human-AI Deliberation. Researchers are encouraged to extend this framework or propose plausible alternatives.
@@ 3.3. Deliberative AI: Design Considerations and Overall Structure
Inspired by human-human deliberation practices and measurements, we derive a set of design considerations for a Deliberative AI. Then, we introduce its overall architecture proposed in accordance with these considerations.
According to deliberative theories (Black et al., 2010; Habermas, 2005; Thompson, 2008) and practices (Lord and Tamvaki, 2013; Steiner et al., 2005), Discourse Quality Index (DQI) (Steenbergen et al., 2003) and its improved versions (Bächtiger et al., 2022, 2009) are the most comprehensive and widely used guideline for assessing human deliberation. We adapt DQI to our AI-assisted decision-making context and summarize the following design considerations (DCs):
@@
Figure 3. The Architecture of Deliberative AI. Our design integrates both a domain-specific model and a Large Language Model, enabling the AI to engage in natural communication with humans while also harnessing domain knowledge derived from the specialized model.
@@
In Sec. 4, we will empirically explore a subset of these dimensions with a specific instantiation of the proposed Human-AI Deliberation framework and Deliberative AI in a graduate admission task.
4. Instantiating the Framework: Graduate Admission Prediction
@@ 4.1. Task, Dataset and AI Model
We choose to use graduate admission as an illustrative task to demonstrate how to instantiate the proposed Human-AI Deliberation framework. In this task, participants decide on admitting or rejecting applicants to a U.S. university based on their profiles. We chose this task for two key reasons. First, this task is widely used in AI-assisted decision-making research (Cheng et al., 2019; Zhang et al., 2023; Echterhoff et al., 2022; Zhao et al., 2023), with real-world universities employing AI algorithms for decision consistency and workload reduction (Waters and Miikkulainen, 2014; Pangburn, 2019). Second, graduate admission often involves deliberation among committee members (Zhang et al., 2023), making it ideal for studying the effects of our proposed Human-AI Deliberation.
@@
To build a domain-specific model (DS-model) that can generate suggestions, we trained a multi-category linear model using a 70% random split of the dataset as in (Cheng et al., 2019). We employed a linear regression model as a decision classifier, discretizing the predicted responses into one of the four decision labels. Consistent with common practices (Echterhoff et al., 2022), we further binarized the original labels, mapping strong/weak reject to “reject” and strong/weak accept to “accept” as the ground truth for assessing AI model’s and participants’ prediction accuracy. The trained model achieved an 80% accuracy on the remaining 30% test set. The task samples used in the study were selected from the test set. It should be noted that in order to explore the effect of Human-AI Deliberation, we specially selected cases from the test data that are challenging for both humans and AI, so that the performance of AI is around 50%. But this does not mean that the AI model used in our study is an unrealistically low-performance AI, just that we focus on difficult task cases.
4.2. Implementation of Human-AI Deliberation
4.2.1. Human-AI Thought Representation and Alignment
@@
In our Human-AI Deliberation framework, the initial step involves both humans and AI externalizing their thoughts. We employ feature contribution (Liao and Varshney, 2021; Liao et al., 2020) to represent their weight of evidence (WoE) along each decision dimension (Alvarez-Melis et al., 2021). Feature contribution is represented by contribution scores indicating the positive or negative influence of each feature $x_{i}$ on the final prediction $y$. In graduate admission tasks, we treat each attribute in an applicant’s profile as a dimension, and feature contribution requires humans and AI to assess the influence of each dimension on the final decision. It provides a common ground for AI and humans to express, compare, and initiate discussions about their thoughts at the feature level (see Table 1).
Table 1. The mapping between AI “thoughts” and human thoughts in the Weight of Evidence framework.
Component AI “thoughts” Human thoughts
Decision Accept or reject one applicant Accept or reject one applicant
Dimension Feature/Attribute (e.g., GPA) Feature/Attribute (e.g., GPA)
Opinion A feature’s contribution (e.g., GPA: + 13%) Weight of evidence (e.g., GPA: -5%)
Evidence Training data Data or other knowledge held by people
For the human side, WoE can be interpreted as the influence of an attribute on the total chance of an applicant being admitted. For the AI side, we used SHAP (SHapley Additive exPlanations) (Lundberg and Lee, 2017), a widely-used explainable AI algorithm, to generate AI feature contributions. SHAP values indicate both the direction and strength of a feature’s impact on predictions. SHAP offers two key advantages: it captures feature interactions, mirroring human decision-making, and its additive nature resembles how humans combine evidence weights for or against options (Baumann et al., 2014). However, when applied in the context of Human-AI Deliberation, two limitations arise: (1) Raw SHAP values can be challenging to interpret directly. To better convey feature influence to non-expert humans, we converted SHAP values into probabilities by adopting a regression model and transforming the four-category label range into a 0-100% range (in this way, the generated SHAP value can be directly mapped to 0-100%)  [^2] Other conversion methods include using the TreeExplainer from the SHAP library (and setting model_output=“probability”) to generate the probabilistic SHAP value. Because the admission prediction used in this paper can be naturally converted into a probability problem, we adopted more direct category label-probability mapping. (2) SHAP values only tell users how important a feature is, but not why it is important. To bridge this gap, we propose generating “meta-explanations” that use Knowledge Extractor (Sec. 3.3) to extract evidence (e.g., data patterns) from training data for each dimension, enhancing AI’s transparency during deliberation.
4.2.2. Implementation of Deliberative AI
@@
I. Communication Layer. This layer serves as a vital bridge between humans and the DS-model, facilitating effective communication by comprehending human inputs and crafting relevant responses.
I-1. Intention Analyzer. We harnessed the language capabilities of LLMs [^3] In this study, we use GPT-3.5-turbo as the language model, which has sufficient conversational power and is freely accessible (relative to GPT-4). We will use “LLM” throughout this paper for consistency. to discern human intentions and targeted dimensions in discussion. To formulate effective prompts for intention analysis, we conducted a pilot study to gather common dialogues around graduate admission decisions, including questions, arguments, critiques, and challenges. In the pilot study, we developed a preliminary version Deliberative AI (with imperfect deliberative discussion capability) to carry out conversations with 30 participants from Prolific444www.prolific.co. Each participant expressed their opinions on various dimensions of applicant profiles and engaged in discussions with the AI, particularly focusing on conflicting opinions. This yielded 226 human deliberative statements. To extract diverse intentions from these statements, two authors conducted qualitative coding using thematic analysis (Hsieh and Shannon, 2005), and the results are summarized in Table 2. We then iteratively refined LLM prompts based on the collected data and built an ”Intention Analyzer” with a 96% accuracy in identifying themes of participant statements. Specific prompts are available in the supplementary materials.
@@
I-2. Deliberation Facilitator. This component addresses DC2 (Justification Rationality) and DC5 (Respect and Agreement), as discussed in Sec 3.3, by designing corresponding LLM prompts. In particular, we instruct LLM to (1) Demonstrate a nuanced understanding of the human’s statement; (2) analyze the specific content of the person’s statement; and (3) provide a thoughtful and critical response. For detailed prompts, please refer to the supplementary materials.
I-3. Argument Evaluator. The main function of this component is to assess the strength of a person’s statement, which informs updates to AI opinions. Drawing from established theories in human argumentation evaluation (Harris, 1997; Van Eemeren et al., 2002; Van Eemeren and Henkemans, 2016), we devised a comprehensive scoring mechanism with nine key items: Clarity, Relevance, Evidence, Logic, Consistency, Counterarguments, Depth, Credibility, and Alignment. These criteria are integrated into a prompt, guiding the LLM to evaluate human statements. We then average and scale the scores to obtain the overall human argument strength $S_{Human}$ (from 0 to 1; 0: weakest, 1: strongest). Additional details, including scoring schemas and prompts, can be found in the supplementary materials.
In summary, the communication layer can engage in general interactions with humans. To imbue it with specific model opinions and knowledge, we require a control layer to mediate between the LLM and the DS-model.
II. Control Layer. This layer manages the querying and extraction of specific DS-model opinions and knowledge while controlling the entire conversation flow.
@@
II-2. Knowledge extractor. Based on the attributes/dimensions and intent types identified by the “Intention Analyzer” (see Table 2), we developed a series of query functions to extract relevant data knowledge from the DS-Model. These functions help pull evidence for the LLM to generate responses in deliberative discussions appropriately. We established a mapping between the recognized intent type and the query function and called different query functions based on the recognized intent type. Below is a brief overview of the designed functions corresponding to different human intent types. Please refer to the supplementary materials for detailed codes and examples.
  • Distribution/Level of an attribute’s value: Function get_distribution (attribute) calculates attribute value percentiles within the applicant pool, along with contextual comparisons (with minimum, maximum, quartiles, mean, and median).
  • Overall attribute importance: Function get_global_feature_importance (attribute) returns global importance. Function get_correlation (attribute) provides Pearson correlation. Function get_influence_on_admission_chance (attribute) calculates admission chance changes for varying attribute values.
  • Contribution of an attribute: Function get_current_value_influence (attribute) calculates admission chance differences when an attribute is randomized.
  • Contrastive Evaluation: Function get_contrastive_admission_chance (attribute, contrastive) computes admission chance differences with a contrastive value.
  • Holistic review of multiple attributes: Function get_holistic_analysis (attributes, fixed_attributes) evaluates attribute impacts in specific scenarios, e.g., GPA percentile in top-ranked schools.
@@
II-4. Opinion Update Controller: We updated the AI’s opinions by taking into consideration: (1) the current opinions of both the human ($O_{Human}$) and the AI ($O_{AI}$) on the discussed attribute, (2) the strength of the human’s argument ($S_{Human}$, see Argument Evaluator), and (3) the AI’s uncertainty ($U_{AI}$) measured and calibrated via Uncertainty Quantification 360 toolbox (Ghosh et al., 2021) (the uncertainty ranges from 0 to 1: the closer to 1, the more uncertain AI’s prediction is). We propose the following formula to update the AI’s opinions ($\hat{O}_{AI}$) on an attribute based on these factors, inspired by result aggregation in crowd intelligence (Lyon and Pacuit, 2013; Galton, 1907):

### v2

@@ 3.1. Integrating Deliberation into Decision-Making
To structure the deliberation process, we break down decision-making problems and human-AI thoughts into four components (Figure 2 (a)): (1) Decision: The overall choice to be made for a specific problem; (2) Dimension: An aspect considered when forming the overall decision. In tabular datasets commonly used in decision-making (Ghai et al., 2021; Wang and Yin, 2021), dimensions often correspond to attributes or sets of related attributes, such as academic excellence or research ability in a graduate admission task; (3) Opinion on a dimension: The assessment of a dimension’s impact on the overall decision (e.g., academic excellence contributes +50% to admission probability); (4) Evidence: The foundation for forming an opinion. For humans, evidence may include facts, heuristics, or personal experiences (Simon, 1990, 1997). For AI, evidence is often rooted in the information encoded in its training data.
We propose deliberation at the dimension level, focusing on how each dimension supports or opposes the overall decision. Both humans and AI must substantiate their opinions with evidence and assess its credibility and probative value (Baumann et al., 2014). This process generates the Weight of Evidence (WoE) (Cartwright and Stegenga, 2011), which quantifies evidence significance/importance relative to alternatives (Peirce, 2014; Good, 1950; Gough, 2007). WoE is widely used in decision-making for its intuitive meaning and practical implementation (Saaty, 2008; Gough, 2007). In the rest of this paper, we will use “opinion” and “WoE” interchangeably.
@@
Figure 2. The architecture of Human-AI Deliberation. (A) Illustrates the Weight of Evidence (WoE) concept in decision-making, showcasing how decision-makers assess evidence across dimensions to shape opinions and arrive at a final decision. (B) Presents the Architecture for Human-AI Deliberation, with key activities (shown in grey boxes) and potential design space (shown in dashed-line boxes).
@@
3.2. An Architecture of Human-AI Deliberation
Based on the WoE-centered decision-making approach, we propose Human-AI Deliberation, an architecture to stimulate deliberative processes between humans and AI (Figure 2). This architecture comprises the following essential activities:
  • Elicitation of Thoughts: Human and AI start with articulating their dimension-level perspectives on the decision problem. While AI presenting its “thoughts” (e.g., in the form of feature importance explanation) is rather common in AI-assisted decision-making (Liao and Varshney, 2021), this activity also encourages individuals to clarify their ideas and examine their reasoning, which prompts analytical thinking in human (Milkman et al., 2009; Buçinca et al., 2021). Two aspects of this activity require careful design. First, AI thought elicitation demands a good balance between human information needs and interpretability (Mishra and Rzeszotarski, 2021; Abdul et al., 2020). Second, human thought elicitation, while encouraging thoughtful reasoning (Milkman et al., 2009; Buçinca et al., 2021), can impose a potential workload. It thus demands suitable, friendly interface designs.
  • Alignment of Human-AI Thoughts: As human’s and AI’s viewpoints and the process they form those viewpoints may diverge (Kaufman and Kirsh, 2022; Miller, 2019; Holzinger et al., 2019), this activity is tasked with establishing a common language for the two parties to compare their WoE and determine the extent of discrepancy. Proper assessment and presentation of human-AI WoE differences can help effectively navigate humans’ attention and efforts in the subsequent activities (Boggust et al., 2022).
  • Discussion: This activity fosters constructive discussions where humans and AI substantiate their opinions, clarify evidence choices, and explain weight assignments. It promotes critical thinking, reduces biases, and highlights differences between parties (Milkman et al., 2009; Larson et al., 1994; Hirokawa, 1985). With a broad design space, it must be tailored to specific decision tasks, considering factors like content, style, leadership (who initiates and leads the conversation), and duration. A potential solution is adapting human-human discussion (Milkman et al., 2009; Larson et al., 1994; Hirokawa, 1985) to the human-AI discussion contexts.
@@ 3.2. An Architecture of Human-AI Deliberation
In summary, the proposed Human-AI Deliberation architecture includes four interlinked activities and requires appropriate designs for both the decision-making interface and the AI (as shown in the dashed box in Figure 2 (b)). Since interface design is task-dependent, we focus on the design of Deliberative AI in the next subsection.
@@ 3.3. Deliberative AI: Design Considerations and Overall Structure
Based on deliberative theories (Black et al., 2010; Habermas, 2005; Thompson, 2008) and practices (Lord and Tamvaki, 2013; Steiner et al., 2005), the Discourse Quality Index (DQI) (Steenbergen et al., 2003) and its improved versions (Bächtiger et al., 2022, 2009) provide a comprehensive framework for assessing human deliberation. We adapt DQI to AI-assisted decision-making and derive the following design considerations (DCs):
@@
Figure 3. The architecture of Deliberative AI which integrates a domain-specific model and a Large Language Model, enabling the AI to engage in natural communication with humans while also harnessing domain knowledge derived from the specialized model.
@@
In the next section, we will provide a detailed description of how we implemented Human-AI Deliberation and Deliberative AI architecture in the context of a graduate admissions task.
4. Instantiating the Architecture: Graduate Admission Prediction
@@ 4.1. Task, Dataset and AI Model
We choose to use graduate admission as an illustrative task to demonstrate how to instantiate the proposed Human-AI Deliberation architecture. In this task, participants decide on admitting or rejecting applicants to a U.S. university based on their profiles. We chose this task for two key reasons. First, this task is widely used in AI-assisted decision-making research (Cheng et al., 2019; Zhang et al., 2023; Echterhoff et al., 2022; Zhao et al., 2023), with real-world universities employing AI algorithms for decision consistency and workload reduction (Waters and Miikkulainen, 2014; Pangburn, 2019). Second, graduate admission often involves deliberation among committee members (Zhang et al., 2023), making it ideal for studying the effects of our proposed Human-AI Deliberation.
@@
To build a domain-specific model (DS-model) that can generate suggestions, we trained a multi-category linear model using a 70% random split of the dataset as in (Cheng et al., 2019). We employed a linear regression model as a decision classifier, discretizing the predicted responses into one of the four decision labels. Consistent with common practices (Echterhoff et al., 2022), we further binarized the original labels, mapping strong/weak reject to “reject” and strong/weak accept to “accept” as the ground truth for assessing AI model’s and participants’ prediction accuracy. The trained model achieved an 80% accuracy on the remaining 30% test set. The task samples used in the study were selected from the test set.
4.2. Human-AI Thought Representation and Alignment
@@
In our Human-AI Deliberation architecture, the initial step involves both humans and AI externalizing their thoughts. We employ feature contribution (Liao and Varshney, 2021; Liao et al., 2020) to represent their weight of evidence (WoE) along each dimension (Alvarez-Melis et al., 2021). Feature contribution is represented by contribution scores indicating the positive or negative influence of each feature $x_{i}$ on the final prediction $y$. In graduate admission tasks, we treat each attribute in an applicant’s profile as a dimension, and feature contribution requires humans and AI to assess the influence of each dimension on the final decision. It provides a common ground for AI and humans to express, compare, and initiate discussions.
For the human side, Weight of Evidence (WoE) represents humans’ perceived influence of an attribute on the overall likelihood of an applicant being admitted. On the AI side, we utilized SHAP (SHapley Additive exPlanations) (Lundberg and Lee, 2017), a widely-used posthoc explainable AI algorithm, to assess feature contribution/importance. SHAP values indicate both the direction and strength of a feature’s impact on predictions. SHAP offers two key advantages: First, it captures feature interactions (e.g., how multiple features jointly influence the final decision), aligning with human decision-making processes. Second, its additive nature mirrors how humans combine evidence for or against options (Baumann et al., 2014). In essence, SHAP allows feature contribution/importance to be linearly aggregate to match the model’s actual prediction no matter the AI model is linear or a complex non-linear neural network.
However, when applied to Human-AI Deliberation, two limitations of SHAP values become evident: (1) Interpretability of Raw SHAP Values: Raw SHAP values (e.g., 2.15) can be difficult for non-experts to interpret directly. To address this, we converted SHAP values into probabilities by using a regression model, mapping the four-category label range to a 0-100% scale. This conversion allows SHAP values to be directly interpreted as probabilities. (2) Explanation of Feature Importance: SHAP values indicate the importance of features but do not explain why a feature is important. To address this, we propose generating “meta-explanations” using the Knowledge Extractor (Sec. 3.3) to extract relevant evidence (e.g., data patterns) from the training data for each dimension. This approach aims to enhance the transparency of AI during the deliberation process.
@@
4.3. Implementation of Deliberative AI
4.3.1. I. Communication Layer.
This layer serves as a vital bridge between humans and the DS-model, facilitating effective communication by comprehending human inputs and crafting relevant responses. Note that all components in this layer are based on our designed prompts, which enable LLMs to play different roles, rather than incorporating additional predictive models.
I-1. Intention Analyzer. We harnessed the language capabilities of LLMs [^3] During our experiment (conducted in August 2023), we used the GPT-3.5 model. For consistency, we will refer to it as “LLM” throughout this paper. to identify human intentions and targeted dimensions in discussion. To formulate effective prompts for intention analysis, we conducted a pilot study to gather common dialogues around graduate admission decisions, including questions, arguments, critiques, and challenges. In the pilot study, we developed a preliminary version Deliberative AI (with basic deliberative discussion capability) to carry out conversations with 30 participants from Prolific444www.prolific.co. Each participant expressed their opinions on various dimensions of applicant profiles and engaged in discussions with the AI. We collected 226 human deliberative statements. To extract various intentions from these statements, two authors performed qualitative coding using thematic analysis (Hsieh and Shannon, 2005), and the results are summarized in Table 1. We then iteratively refined LLM prompts based on the collected data and built an “Intention Analyzer” with a 96% accuracy in identifying themes of participant statements. Specific prompts are available in the supplementary materials.
@@ 4.3. Implementation of Deliberative AI
I-2. Deliberation Facilitator. This component addresses DC2 (Justification Rationality) and DC5 (Respect and Agreement) by designing corresponding LLM prompts. In particular, we instruct LLM to (1) Demonstrate a nuanced understanding of the human’s statement; (2) analyze the specific content of the person’s statement; and (3) provide a thoughtful and critical response. For detailed prompts, please refer to the supplementary materials.
I-3. Argument Evaluator. The main function of this component is to assess the strength of a person’s statement/argument, which later informs updates to AI opinions. Drawing from established theories in human argumentation evaluation (Harris, 1997; Van Eemeren et al., 2002; Van Eemeren and Henkemans, 2016), we devised a comprehensive scoring mechanism with nine key items: Clarity, Relevance, Evidence, Logic, Consistency, Counterarguments, Depth, Credibility, and Alignment. These criteria are integrated into a prompt, guiding the LLM to evaluate human statements. We then average and scale the scores to obtain the overall human argument strength $S_{Human}$ (from 0 to 1; 0: weakest, 1: strongest). We conducted a pilot study to evaluate the reliability of the LLM in scoring human arguments, using data collected from our Intention Analyzer pilot study (Sec. 4.3.1 I-1). Two authors independently scored the arguments using predefined schemas, resolving disagreements to reach consensus. We then calculated Cohen’s Kappa ($\kappa$) to measure agreement between the LLM and human scores. The resulting $\kappa$ value of 0.78 indicates substantial agreement, suggesting that the LLM provides reliable annotations with minimal disagreement in the context of our graduate admission task. Additional details, including scoring schemas and prompts, can be found in the supplementary materials.
In summary, the communication layer enables general interactions with humans. To integrate it with the DS-model’s predictions and knowledge, a control layer is required to connect the two.
4.3.2. II. Control Layer.
This layer manages the querying and extraction of specific DS-model opinions and knowledge while controlling the entire conversation flow.
@@
II-2. Knowledge Extractor. Based on the attributes/dimensions and intent types identified by the “Intention Analyzer” (see Table 1), we developed a series of query functions to extract relevant data knowledge from the DS-Model. These functions help pull evidence for the LLM to generate responses in deliberative discussions appropriately. We established a mapping between the recognized intent type and the query function and called different query functions based on the recognized intent type. In practice, for a decision-making task, an existing dataset is typically available for training the AI model. This same dataset can be used to extract data patterns based on task-specific features, such as the percentile of a single feature value or the combination of multiple features. We recommend that researchers interested in applying this approach to other tasks first identify the data patterns that users are likely to find relevant, and then design the corresponding extraction functions. Below is a brief overview of the designed functions corresponding to different human intent types. Please refer to the supplementary materials for detailed codes and examples.
  • Distribution/Level of an attribute’s value:
    • Function get_distribution(attr_val) calculates attribute value percentiles within the applicant pool, along with contextual comparisons (with minimum, maximum, quartiles, mean, and median).
  • Overall attribute importance:
    • Function get_global_feature_importance(attr) returns global importance.
    • Function get_correlation(attr) provides Pearson correlation between the selected attribute and the admission chance.
    • Function get_influence_on_admission_chance(attr) calculates admission chance changes for varying attribute values.
  • Contribution of an attribute:
    • Function get_current_value_influence(attr) calculates admission chance differences when the value of a selected attribute is randomized.
  • Contrastive Evaluation:
    • Function get_contrastive(attr, contrast) computes admission chance differences compared to a contrastive value.
  • Holistic review of multiple attributes:
    • Function get_holistic_analysis(attr, fix_attr) evaluates the impacts of attribute(s) in specific scenarios, e.g., the impact of the GPA percentile in [top-ranked universities].
@@
II-4. Opinion Update Controller: We updated the AI’s opinions by taking into consideration: (1) the current opinions of both the human ($O_{Human}$) and the AI ($O_{AI}$) on the discussed attribute, (2) the strength of the human’s argument ($S_{Human}$, see Argument Evaluator), and (3) the AI’s uncertainty ($U_{AI}$) measured and calibrated via Uncertainty Quantification 360 toolbox (Ghosh et al., 2021) (the uncertainty ranges from 0 to 1: the closer to 1, the more uncertain AI’s prediction is). We propose the following formula to update the AI’s opinions ($\hat{O}_{AI}$) on an attribute based on these factors, inspired by result aggregation in crowd intelligence (Lyon and Pacuit, 2013; Galton, 1907). It is important to clarify that the term “update” here does not refer to modifications to the domain-specific model itself, such as retraining or fine-tuning. Instead, it pertains to the adjustment of the AI’s expressed viewpoints regarding the current task case. Notably, LLM is not directly involved in updating the domain-specific model’s viewpoints. Rather, LLM generates a strength score for the human’s arguments, with the update being performed using Eq. 1.